The mark scored by Henry in his English ,Maths and Science test were in the ratio 4:5:6.He scored 42 mark less in science than the total mark of his English and maths test. How many marks did he get for his English test??
Answer by ikleyn(52944)

.

The marks are  4x (English),  5x (Math)  and  6x (Science).


The equation is


    6x = 4x + 5x - 42


Simplify and solve


    6x = 9x - 42

    42 = 9x - 6x

    42 = 3x

     x = 42/3 = 14.


ANSWER.  The mark in English is  4*14 = 56.

Solved.